OpenStudy (anonymous):

solve 2sinxcosx=0 sin2x=0 2x=sin^-1 (0) Is the soln. x=pi/n or x=pi n/2?

OpenStudy (anonymous):

oh, there's a typo... x=pi n or x=pi n /2

OpenStudy (tkhunny):

You an do it without alteration. sin(x) = 0 at x = 0 and pi cos(x) = 0 at x = pi/2 and 3pi/2 Or, you can use the transformation you selected. sin(2x) = 0 at 2x = 0, pi, 2pi, 3pi giving x = 0, pi/2, pi, 3pi/2 Isn't it interesting that mathematics works?!

OpenStudy (anonymous):

so... you're saying that the latter, x=pi n /2 is the correct answer?

OpenStudy (anonymous):

\[0,\frac{\pi}{2},\pi,\frac{3\pi}{2},2\pi,...\] any place where either sine is zero or cosine is zero

OpenStudy (anonymous):

if you want to abbreviate this you can write \[x=\frac{n\pi}{2},n\in \mathbb{Z}\]

OpenStudy (anonymous):

oh... now I see what you were trying to tell me. hahaha but representing it would be pi n /2, correct?

OpenStudy (anonymous):

oh, ok thank you so much! my firend gave me a wrong answer LOL

OpenStudy (anonymous):

don't trust friends

OpenStudy (anonymous):

:) hahaha ok i wont

OpenStudy (anonymous):

lesson learned? :)

OpenStudy (anonymous):

u r one of my favorite ppl on here~ hahaha keep at it!